# Information Seeking Using AI Assistants

### v1

@@ IV-A Information Seeking with AI tools (RQ1)
IV-B Impact on Perceived Productivity (RQ2)
Most of our participants reported the use of AI tools for information seeking having a positive impact on their productivity. More specifically, participants cited time savings when seeking information for trivial, routine, or repetitive tasks. Most notably, AI tools support quicker finding and consolidation of relevant information and sometimes offer step-by-step guidance on how to make use of the information. Participants also mentioned AI tools reducing time spent blocked by obstacles, as AI tools offer quick solutions and guidance, allowing for efficient problem-solving and smoother project progression. For some participants, the time savings in these more trivial tasks also supports their ability to focus on more complex tasks.
“I use an AI assistive tool [when] staring at a blank screen and trying to figure out, okay, what is it that I need to do next? What component do I need to do? What do I need to start with right now?… One of the [AI] tools will break that down into smaller steps and help break the flow a bit to help me get back into the flow.”[P10]
In fact, despite the potential for new tooling to be disruptive to developer workflows, which can have a negative impact on productivity, participants from our survey and interviews indicated AI tools may not be as counter-productive when information seeking. One participant emphasized this distinction, stating, “The efficiency gained is worth the loss of flow. But for me, flow is less about thinking and more about implementing. The AI is doing the implementing, so the work I am doing is overcoming the roadblocks, not doing the easy work that flows.”
For most of our survey respondents, AI tools at the very least have no impact on their workflow (109), where respondents reported that their workflow was about the same when using AI tools as when they do not. For some respondents, mostly those with fewer years of development experience, using AI tools for information seeking actually improves their flow (78). Those who felt AI tools improved their flow, the rationale was most often not having to look at as many sources, being provided step-by-step guidance, and faster learning. As indicated by many participants, using AI tools supports getting “a lot of stuff in one place” rather than “searching multiple links” returned from, for example, from a Google search.
@@ IV-B Impact on Perceived Productivity (RQ2)
While time savings was the most common metric of productivity among our participants, our efforts uncovered other ways developers think about productivity when using AI tools for information seeking. When considering time savings, this can be the time required to complete tasks or projects, debug and resolve issues, and make decisions. We also found that output quantity was another metric; this refers to the number of tickets or tasks completed per day or sprint, as well as publication output for research and development teams. One participant noted that they could “get a ticket that has been scoped for several days or a week done in a half an hour”. Another stated when they can close “[more] tickets [than usual] in a day” or “[multiple] feature development and a couple of bug fixes” then they know “the tool is assisting [me] in the right way.” Some participants also discussed other considerations that affect how they measure their productivity when using AI tools. For example, some participants emphasized the role of solution quality in their productivity, which includes things like frequency of issues arising and the number of code review iterations required.
We found that some productivity-related considerations had an impact on the decision to use an AI tool for information seeking to begin with. One such consideration is cost-effectiveness which is measured by the resources required (e.g., number of developers involved). As put by one participant, “ I would say, if we are able to finish this chat analysis project. Oh, my God! That would save thousands of person hours.”
@@
Our participants also highlighted the impact of using AI tools for information seeking beyond the individual developer. For some, there have been visible impacts to the projects they are working on when using AI tools. Participants cited AI assistants making it easier to parse large volumes of documentation to help pinpoint and resolve issues faster and providing code optimization strategies that speed up application. Some also discussed AI tools helping them clarify requirements and consider edge cases, leading to faster sprint completion times and more robust solutions. When discussing their own experiences, one participant described a concrete example of this kind of support:
“I needed to handle a bulk of requests from customers’ mobile devices. Initially, I assumed that the number of requests would not exceed a certain limit in a real-life scenario. However, there was an instance where this assumption was proven wrong, leading to a significant surge in requests. I quickly consulted with GPT, and we implemented a quick fix to resolve the issue that day. Since then, I’ve started considering potential edge cases more rigorously. I now regularly check with GPT, asking for potential edge cases based on my approach. The suggestions I receive help me anticipate potential issues, allowing me to proactively incorporate necessary adjustments into my design to avoid future problems.”[P6]
On the flip side, participants also discussed the negative impacts AI tool use can have. Most noticeable is the impact AI tools have on software teams. Our participants report reduced face-to-face interactions and organic knowledge sharing, which can impact team cohesion and collaboration. As described by one participant, some may be more inclined to use AI even when collaborators are an option:
“I would say one negative effect of it has been that we don’t have as much synergy with each other. And I think this also ties in sometimes to the fact that we’ve mainly gone to working remote. Because we’re no longer forced to bounce things off each other. I would say we’re less incentivized to bounce ideas off each other, ask for help on concepts, and I believe some of that organic conversation has gone missing. So for me … I happen to work with a very renowned scientist in that area. But, you know, in my mind, I’m more likely to go to like AI for help.”[P11]
Participants feel especially more comfortable using AI tools when they are less knowledgeable about something, as they may feel self-conscious about approaching someone more expert.
“We used to have different Slack channels for like Python help, or Excel, or web development. Those have kind of, I wouldn’t say gone silent, but they’re, if anything, mainly used for jokingly sharing a really bad answer from an AI. But at the end of the day I think it’s preferable for them to use AI instead of outing themselves not knowing a certain coding convention or making a really silly error, because it’s not anonymized on Slack.” [P11]
IV-C Impact on Skill Development (RQ3)
@@
Many participants discussed how using AI tools when seeking information supports their learning and skill-building. Most often, participants mentioned the effectiveness of AI tools for filling in knowledge gaps and providing practice problems that support hands-on learning. For these participants, using AI tools in this way helps reinforce new concepts and improve problem-solving skills. In fact, the most frequently cited use for AI tools was explaining and clarifying concepts relevant to their tasks or goals. Participants often discussed an appreciation for AI tools’ ability to break down complex topics into more understandable parts. Our participants also reported frequently using AI tools for learning best practices, citing them as valuable resources for understanding industry standards and improving coding techniques.
AI-driven information seeking can also support the discovery and adoption of new solutions. An overwhelming majority of survey respondents reported using AI for learning new technologies (119). Participants in our study found AI tools useful for exploring innovative approaches and alternative methods for solving programming challenges, where they noted being able to get acquainted with new technologies without the overhead of seeking and learning about each individually. One participant noted being “more willing to use new tooling I don’t understand, barely understand, or rarely work with.” AI tools have also played a role in building other participants’ confidence in using new libraries or technologies, though some still reported discomfort in applying the knowledge gained from AI tools independently to practical scenarios. For one participant, this can be a product of overreliance, stating, “if you keep looking at ChatGPT for more and more alternate solutions, it just makes you lose your confidence.”
While many of our interviews suggest some usefulness of AI tools for learning new things, a high proportion of respondents from our survey indicated that they experience a higher learning curve when learning about and understanding new technologies using AI. In fact, for some participants, AI tools are not suited for supporting their learning, despite their ability to provide personalized responses. For some, the information presented by AI tools often provide inadequate levels of detail. As stated by one participant:
@@
When discussing the ways in which they feel AI tools impact their skill-building, many participants cited appreciating the ability to expand their knowledge and technical skillset more quickly with AI tools. However, there were also numerous concerns surrounding the skill building aspect of using AI tools for meeting information needs. From our survey, we found that most of our respondents have concerns regarding using AI tools for learning and skill-building because use in this context could lead to knowledge gaps (88), overreliance on AI tools (94), and poor problem solving skills (87).
@@
One notable drawback that emerged from our interviews is the decrease in the ability to build creative problem-solving skills when engaging AI tools for meeting information needs. As indicated by participants, this can lead to more fundamental issues such as decreased ability to code from scratch and reduced understanding of the technologies being used.
@@
“It’s like when we were doing calculations on our own. But then calculators came in. So we don’t calculate it in our mind anymore Whenever I come across a problem I just put it into GPT. I just don’t think about it on my own anymore, because I am relying a lot on AI tools… I just put it in there if I think it will help. It has greatly reduced my [ability to think independently].”[P7]
@@
This includes diminished learning regarding use of traditionally fundamental software development tools and techniques such as command line, scripting, and debugging, where they can now “just throw the error, or whatever problem,…to ChatGPT and it will fix it.” This is exaggerated by the fact that the knowledge provided by AI tools is often limited to the specific question asked, often lacking the depth and breadth required for comprehensive understanding. In addition, iterative interactions are often required where the user must continuously refine their questions to receive useful answers. Some participants emphasized the challenge of not knowing what to ask or how to ask AI tools for information when learning about something new. As one participant noted, “it only covers the range of questions I have in my mind…it is like giving an answer to your question, but not delivering a session on one topic.”
While using AI tools can reduce creativity and critical thinking, some participants in our study also noted that AI-assisted information seeking can formulate a different kind of creative and critical thinking skillset. For example, according to some participants, correcting or refining an AI tool’s suggestions requires creativity, critical thinking, and problem solving skills. One participant made this notion concrete, stating:
@@
“…and when I’m done I approach GPT for practice. So I’ll actually ask the GPT to give me [practice] problems that are, for example, I’ll take a component and I will ask GPT to give me [problems to solve] in different scenarios. So then I would … start thinking of [how to solve] them. And then I would ask it again to give me the solution when I’m done [figuring out] my own so that I can compare and think. It helps me think in a different way.”[P2]
@@
This process can also foster creativity by sparking new ideas and approaches the developer may not have initially considered. A handful of participants reported that AI tools actually improve creativity, as they provide a foundation upon which developers can build and expand their ideas. As noted by one participant, “I think it promotes creativity. Because you may be introduced to new ways or new patterns of doing something…I think that it sometimes, by being too wordy, will like accidentally give people ideas that they otherwise wouldn’t have, or like expose people to ideas that they otherwise wouldn’t have. Yeah, I’d say it’s like a net positive win in the creativity department”[P17].
@@
Our survey respondents also indicated that they use AI tools for both learning about (114) and integrating new technologies (113) in their work. However, from our interviews, we found there may be differences in the impact of using AI tools for learning about new technologies rather than integrating new technologies. One participant summarized this well, stating:
@@ V Discussion
V-A AI Tools Cannot (Always) Replace Developers
Two major, and related, concerns that have emerged with the rapid advancement and integration of AI tools in software development is that eventually nobody will have to learn to program anymore  [^2] https://www.techradar.com/pro/nvidia-ceo-predicts-the-death-of-coding-jensen-huang-says-ai-will-do-the-work-so-kids-dont-need-to-learn and AI will replace developers [26]. These concerns are likely heightened by the development of tools like ChatGPT and Gemini that provide human-like interactions. Furthermore, as our findings suggest developers may be engaging with each other less frequently when using AI tools, this raises important concerns regarding the future of collaborative software development; if developers are hyper-focused on AI tools facilitating their work, perhaps because they are too intimidated to talk to a teammate, this could greatly impact the ability for teams to coordinate heavier tasks, like integration, and to identify when teammates are duplicating efforts. While our findings suggest that there may be tasks that AI tools can perform more efficiently, and even effectively, than developers, they also emphasize how unlikely this future is.
Prior work has already emphasized the importance of social support in programming [27] and provided insights into the distinctions between AI and developers [28]. Our findings provide further evidence to debunk these myths by emphasizing the role of the developer in AI tools’ ability to support information seeking and meet their needs. For example, one theme throughout our findings was that an AI tool is only as effective as the interactions the developer has with it. This is the whole premise of efforts in prompt engineering that aim to understand how to get the best outcomes from AI tools [29, 30, 31]. But most importantly, our work amplifies how far AI tools are from being self-sufficient, relying heavily on developer knowledge and expertise for guidance towards relevant, high-quality information that can meet their goals and facilitate efficient task completion. This suggests that developers may be more often perceiving these tools as copilots that work in synergy with them rather than independently automating a handful of tasks, such as generating boilerplate code. Even then, according to our findings, there are some gaps that AI tools currently do not and can maybe never fill, such as trust, personality, and focused expertise, that make software development both fruitful and enjoyable.
V-B AI-Assisted Info Seeking Is Productive (Sometimes)
Another concern that has been raised in the software engineering community is regarding the negative impacts that AI tools can have on productivity and the ability to measure the impact it is having [12, 32, 33, 2]. Our findings suggest that developers perceive that using AI tools for information seeking can lead to significant and worthwhile productivity gains. In fact, despite any potential disruption to existing workflows when using AI tools, and the impact prior work suggests this may have [15, 34, 20], our findings suggest that the efficiency gains may be worth the loss of flow or even reduce the disruptions by reducing context switches that are often inherent in seeking information [35, 36]. In fact, we found that there are a range of scenarios where developer believe AI-assisted tools are increasing their productivity. Prior work has already emphasized the potential productivity gains from using AI-assisted tools for code completion [12]. Our findings suggest these potential productivity gains extend much beyond code completion into tasks such as consolidating large volumes of information, providing step-by-step guidance, filling in knowledge gaps, and dealing with the “blank page problem” of not knowing how to get started.  [^3] https://www.forbes.com/sites/forbesagencycouncil/2023/07/20/generative-ai-and-solving-the-blank-page-problem/
Our efforts also provide novel insights into the specific productivity-related considerations when measuring productivity and impact of AI tool use in information seeking. The most dominant theme across our findings was that of simplicity. As found in prior work focused on using AI tools for coding tasks, we found that information seeking with AI tools provides the biggest boost in productivity when the ask is simple. In the context of meeting information needs, this means avoiding questions that requires knowledge that cannot be easily accessed. For example, an AI tool can easily curate relevant content for a given tool, technology, or concept, which would save the time it would have taken for a developer to do it themselves. However, asking the AI to provide information to resolve a specific issue or complete a task can lead to wasted time, as well as less than favorable outcomes, given there may be context or insight not being considered by the AI.
V-C AI Tools Cannot (Alone) Support Skill Development
As suggested by prior efforts, and emphasized in our findings, AI tools are increasingly being adopted to increase developer efficiency when learning [37]. Our work provides detailed insights into the potential for AI tools to increase efficiency when learning and developing skills, which is how developers advance from novice to expert as programmers and in their workplace [10]. Research has investigated the ability to learn using AI tools [38, 39, 40]. However, most of this work has been situated in classroom settings where the methods and goals for learning are different than for developers, as our findings also emphasize that they most often acquire additional or new knowledge through the completion of their work tasks.
This important distinction points to a need to better understand and facilitate responsible use of AI tools in the context of building expertise. This is especially the case given the open-ended nature of most AI tools, where they can facilitate both task-oriented and goal-oriented information seeking. In our efforts, we found that AI tools are best suited for supporting the retrieval and consolidation of information, which can indirectly facilitate or potentially expedite learning or skill building. However, our findings also emphasize the importance of foundational knowledge in the effectiveness of AI tools in this context and the risks of over-reliance. This suggests a necessary shift in how and when AI tools should be used, promoting the use of AI tools when learning or understanding is the goal (e.g., what are the best practices for X?) within the context of a task (e.g., fix this bug) as opposed to using them to complete the task for them. In fact, our findings suggest if the goal becomes too close to the task (e.g., is this fix sufficient?), the nature of AI tools’ communication style can lead to an incorrect understanding of solution quality.
This also indicates a shift in the perception of important skills for developers. For example, many participants in our study felt it is becoming less important to exhaustively learn programming languages and more important to learn the fundamentals and best practices, as only when the necessary foundations are present can AI tools effectively and efficiently fill in the gaps. This also supports the ability to more effectively utilize AI tools as learning aids, as our findings indicate it may be difficult to use AI tools in this way without having the knowledge required for adequate prompting and iteration. These considerations are especially important given the impact our findings suggest appropriate (or inappropriate) use of AI tools can have on developer confidence in their knowledge and skills (and ability to apply them in practice).
Our findings also point to the potential for a future shift in where knowledge is found, stored, and accessed and the impact that may have on AI-assisted tools. Evidence of this potential shift can be found in prior work [41] as well as trends over the past couple of years on the StackOverflow Developer Survey, where the percentage of developers who frequently visit StackOverflow decreased from 2022 [42] to 2024 [43]. Our findings suggest this decrease may be happening due to an increase in use of AI tools for goals that developers would typically visit StackOverflow to fulfill. While on the surface this may not appear problematic, reduced interactions on these platforms will inevitably lead to fewer documented question and answer text which impact the ability for AI tools to learn from relevant, up-to-date information. This could also become an issue for StackOverflow and other knowledge repositories where the ability to scale becomes virtually non-existent.
VI Threats to Validity

### v2

@@ 4.1. Information Seeking with AI tools (RQ1)
4.2. Impact on Perceived Productivity (RQ2)
Despite the potential for new tooling to be disruptive to developer workflows, participants from our survey and interviews indicated AI tools may not be as counter-productive when information seeking. One participant emphasized this distinction, stating, ”The efficiency gained is worth the loss of flow. But for me, flow is less about thinking and more about implementing. The AI is doing the implementing, so the work I am doing is overcoming the roadblocks, not doing the easy work that flows.” For most of our survey respondents, AI tools at the very least have no impact on their workflow (109), where respondents reported that their workflow was about the same when using AI tools as when they do not. For some respondents, mostly those with fewer years of development experience, using AI tools for information seeking actually improves their flow (78). For those who felt AI tools improved their flow, the rationale was most often not having to look at as many sources, being provided step-by-step guidance, and accelerated learning.
@@ 4.2. Impact on Perceived Productivity (RQ2)
To better contextualize the perceived productivity reported by developers, we asked participants to describe what productivity meant to them and how they measured it. While time savings was the most common metric of productivity among our participants, our efforts uncovered other ways developers think about productivity. When considering time savings, this can be the time required to complete tasks or projects, debug and resolve issues, and make decisions. We also found that output quantity was another metric; this refers to the number of tickets or tasks completed per day or sprint, as well as publication output for research and development teams. One participant noted that they could “get a ticket that has been scoped for several days or a week done in a half an hour”. Another stated when they can close “[more] tickets [than usual] in a day” or “[multiple] feature development and a couple of bug fixes” then they know “the tool is assisting [me] in the right way.”
Some participants also emphasized the role of solution quality in their productivity, which includes things like frequency of issues arising and the number of code review iterations required.
@@
Our participants highlighted the impact of using AI tools for information seeking beyond the individual developer. For some, there have been visible impacts to the projects they are working on when using AI tools. Participants cited AI assistants making it easier to parse large volumes of documentation to help pinpoint and resolve issues faster and providing code optimization strategies that speed up application. Some also discussed AI tools helping them clarify requirements and consider edge cases, leading to faster sprint completion times and more robust solutions. When discussing their own experiences, one participant described:
”I needed to handle a bulk of requests from customers’ mobile devices. Initially, I assumed the number of requests wouldn’t exceed a certain limit, but there was an instance where this assumption was proven wrong, leading to a significant surge. I consulted with GPT, and we implemented a quick fix that day… Since then, I regularly check with GPT for potential edge cases, which helps me anticipate issues and incorporate necessary adjustments proactively.” [P6]
On the flip side, participants discussed negative impacts AI tool use can have, particularly on software teams. Our participants report reduced face-to-face interactions and organic knowledge sharing, which can impact team cohesion and collaboration:
”We don’t have as much synergy with each other… we’re less incentivized to bounce ideas off each other or ask for help on concepts, and that organic conversation has gone missing. I work with a renowned scientist in that area, but I’m more likely to go to AI for help.” [P11]
Participants feel especially more comfortable using AI tools when they are less knowledgeable about something, as they may feel self-conscious about approaching someone more expert:
”We used to have different Slack channels for Python help, Excel, or web development. Those have mainly gone silent except for sharing bad AI answers. People prefer using AI instead of outing themselves for not knowing certain coding conventions or making silly errors.” [P11]
4.3. Impact on Skill Development (RQ3)
@@
Many participants discussed how using AI tools when seeking information supports their learning and skill-building. Most often, participants mentioned the effectiveness of AI tools for filling in knowledge gaps and providing practice problems that support hands-on learning. For these participants, using AI tools in this way helps reinforce new concepts and improve problem-solving skills.
Participants also frequently used AI tools for learning best practices, citing them as valuable resources for understanding industry standards and improving coding techniques.
AI-driven information seeking can also support the discovery and adoption of new solutions. An overwhelming majority of survey respondents, including more experienced developers, reported using AI for learning new technologies (119).
Participants in our study found AI tools useful for exploring innovative approaches and alternative methods for solving programming challenges, where they noted being able to get acquainted with new technologies without the overhead of seeking and learning about each individually. One participant noted being “more willing to use new tooling I don’t understand, barely understand, or rarely work with.” AI tools have also played a role in building other participants’ confidence in using new libraries or technologies, though some still reported discomfort in applying the knowledge gained from AI tools independently to practical scenarios. For one participant, this can be a product of overreliance, stating, “if you keep looking at ChatGPT for more and more alternate solutions, it just makes you lose your confidence.”
While many interviews suggest usefulness of AI tools for learning, a high proportion of survey respondents indicated they experience a higher learning curve when learning about and understanding new technologies using AI. For some participants, AI tools are not suited for supporting their learning, despite their ability to provide personalized responses, as they often provide inadequate levels of detail. As stated by one participant:
@@
When discussing the ways in which they feel AI tools impact their skill-building, many participants cited appreciating the ability to expand their knowledge and technical skillset more quickly with AI tools, but expressed concerns about potential knowledge gaps (88), overreliance (94), and poor problem-solving skills (87) resulting from AI use for learning.
@@
One notable drawback that emerged from our interviews was the decrease in creative problem-solving abilities, leading to fundamental issues like reduced ability to code independently and diminished understanding of technologies.
@@
“It’s like when we were doing calculations on our own. But then calculators came in. So we don’t calculate it in our mind anymore… I just don’t think about it on my own anymore, because I am relying a lot on AI tools… It has greatly reduced my [ability to think independently].”[P7]
@@
This extends to diminished learning of fundamental development skills like command line and debugging, where developers “just throw the error, or whatever problem,…to ChatGPT and it will fix it.” This problem is compounded by AI providing answers limited to specific questions asked, lacking comprehensive context. Some participants noted the challenge of not knowing how to formulate effective questions when learning something new: “it only covers the range of questions I have in my mind…it is like giving an answer to your question, but not delivering a session on one topic.”
Interestingly, some participants reported that AI fostered a different kind of creative and critical thinking skillset. They described how correcting AI suggestions requires creative problem-solving:
@@
“…when I’m done I approach GPT for practice… I’ll ask GPT to give me [practice] problems… I would start thinking of [how to solve] them. And then I would ask it again to give me the solution when I’m done [figuring out] my own so that I can compare and think. It helps me think in a different way.”[P2]
@@
Some even reported AI improving creativity by exposing them to new approaches: “I think it promotes creativity. Because you may be introduced to new ways or new patterns of doing something…I think that it sometimes, by being too wordy, will like accidentally give people ideas that they otherwise wouldn’t have… Yeah, I’d say it’s like a net positive win in the creativity department”[P17].
@@
Survey respondents reported using AI tools for both learning about (114) and integrating (113) new technologies. However, from our interviews, we found there may be differences in the impact of using AI tools for learning about new technologies rather than integrating new technologies:
@@ 5. Discussion
5.1. The Evolution of Developer Information Seeking
Our research reveals a shift from goal-oriented to task-oriented learning approaches in AI-assisted development. While traditional goal-oriented learning emphasized building comprehensive understanding by exploring available resources in-depth, the current task-oriented paradigm focuses on immediate problem resolution using AI tools as on-demand information providers. This represents a significant shift from established patterns where developers built expertise through reading documentation and social learning (such as peer interactions and collaborative forums) before implementation (Brandt et al., 2009; Zamiri and Esmaeili, 2024).
Our findings suggest that the timing and nature of information seeking may be shifting as well. Developers now typically proceed directly to implementation with AI support, seeking contextual information reactively when encountering specific obstacles. This just-in-time pattern is a contrast from traditional approaches where developers first built foundational knowledge through documentation and peer discussions. While this new approach accelerates immediate task completion, it introduces additional challenges. Another challenge lies in the validation of AI-generated solutions. Unlike traditional information sources that improve through community vetting and peer review, AI-generated content requires individual validation for each instance. This verification burden partially offsets the efficiency gains offered by AI assistance and shifts the responsibility of verification from the community to individual developers. We discuss potential directions for addressing these emerging challenges through developer tool support in Section 5.3
5.2. Developer Productivity and Learning Trade-offs
Developers traditionally build expertise through deliberate practice, pattern recognition, and incremental learning – all of which are well established principles from educational psychology (Campitelli and Gobet, 2011; Cowan et al., 2004; Bor and Seth, 2012; Atkinson and Shiffrin, 1968; Roediger and McDermott, 1995). Self-explanation, reflection, and social learning through mentorship and collaboration have also been crucial components in knowledge transfer and skill development (Boguslawski et al., 2024).
Our research suggests significant trade-offs between immediate productivity gains and long-term expertise building in AI-assisted task completion. While our participants reported, and prior work suggests (Ziegler et al., 2022; Peng et al., 2023; Tung, 2024; Coutinho et al., 2024), AI tools demonstrably increase short-term productivity through faster task completion and reduced blocked time, they also discussed the potential for leveraging AI tools for information seeking to simultaneously impede certain aspects of learning and skill development. We also found that this can lead to decreased confidence in their ability to work with new tools or technologies without the support of AI assistants.
In the current landscape, access to AI-generated solutions is widespread which our findings suggest can reduce the struggle of being productive while learning. However, it can also potentially limit the development of key skills like problem solving and deeper understanding of technical concepts. Furthermore, while AI tools make information more readily available, the fragmented nature of AI-assisted learning may impede the integration of the low-level knowledge acquired into a coherent mental model of software systems and concepts that can be translated to other scenarios (Miller and Read, 2013). This creates complex trade-offs between immediate efficiency and sustained expertise development that warrant careful consideration as AI tools become increasingly integrated into development workflows.
5.3. Directions for AI-Assisted Developer Tools
Central to evolving as an engineer is the ability to build and retain expertise in technical concepts. While our findings emphasize potential trade-offs that may come with using AI-assisted tools for development tasks, they also suggest potential directions for next-generation AI-assisted development tools that better balance immediate assistance with long-term learning support:
Retrieval Augmented Development Environments: While current AI tools commonly used by developers often operate in isolation from project contexts, Retrieval-Augmented Generation (RAG) systems are a promising direction for improving both productivity and solution quality (Lewis et al., 2020). By grounding AI assistance in project documentation, codebase history, relevant forum discussions and technical specifications, RAG systems could help developers more effectively consolidate, leverage, and connect existing knowledge. This could be particularly valuable for tasks like navigating complex documentation, understanding API usage patterns, or debugging specific error cases that require synthesizing information from multiple sources. They could also help automate validation by connecting AI suggestions to trusted sources and providing clear provenance for generated solutions.
Adaptive Learning Systems: Tools need to evolve beyond static assistance to support different learning stages and developer growth. Prior work suggests that adaptive learning systems can be a useful model to replicate in the context of improving developer tools (Johnson et al., 2015). By creating more adaptive AI-assisted developer tools, these systems can better support learning and expertise building by recognizing a developer’s progression from novice to expert and adjusting the depth and style of assistance accordingly. For instance, they might provide more comprehensive explanations for newcomers while offering more context-specific, advanced suggestions as expertise grows.
Task-Specific Fine-tuning: Rather than relying on general-purpose models, AI-assisted tools being used in the context of software development should be specifically fine-tuned for distinct software engineering tasks. This includes specialized models for code review, architectural decision support, security analysis, and learning assistance where each model is optimized for its specific use case and incorporates the most relevant best practices and knowledge.
Knowledge Integration and Sharing: As emphasized in prior efforts, software engineering is a collaborative activity where learning and expertise-building is often facilitated by peers (Boguslawski et al., 2024). To avoid AI-assisted tools replacing or interfering with knowledge sharing among teams and the broader community, these tools could actively facilitate and incentivize knowledge sharing and validation with others. This could include features for documenting AI-assisted solutions, sharing verified responses, and building team-specific knowledge bases that combine traditional documentation with validated AI-generated content.
6. Threats to Validity